OpenStudy (anonymous):

\[v_0=0\] \[x_f-x_i=.074m\]

OpenStudy (anonymous):

wait what ar eyou trying to find

OpenStudy (anonymous):

\[\Delta S=s_f-s_0\] if you let \[s_0= 0\] \[\Delta S=s_f-0=s_f=.074m\] \[S=\frac{1}{2}at^2+0+0\] \[S=\frac{1}{2}at^2\] \[.074=\frac{1}{2}at^2\] \[V_x=at+0\] do you know the velocity at the final distance ... we have 3 unknowns atm, a, t and v

OpenStudy (anonymous):

now it's possible to find this if you're saying that x is actually the y direction so say you have an object being dropped